On January 1, Year 1, Fukisan purchased a new piece of equipment for specialized-furniture manufacturing at a cost of $300,000, inclusive of shipping and installation. At the time of purchase, the equipment had an estimated useful life of 15 years and an expected salvage value of $10,000 at the end of the 15 years.

For future budgeting purposes, Eric Anderson, CFO of Fukisan Inc. has asked you to perform the depreciation expense calculations for Year 2, Year 3, and Year 4.

Use the information above to complete the tasks below. Select the appropriate response from the list.

Using the sum of the years' digits (SYD) method, determine the amount of depreciation expense in Year 4.
Using the double-declining balance (DDB) method, determine the amount of depreciation expense in Year 2.
Assume that the straight-line method is used in Years 1 and 2. At the beginning of Year 3, the estimated remaining useful life is revised to 10 years and the expected salvage value is revised to $5,000. Using the straight-line (SL) method, determine the amount of depreciation expense in Year 3.
Answer #1
Straight line Doble declining Sum of Year Digits
Year 1 $300,000-10,000/15 = $19,333 $300,000*13.33% = $39,990 $300,000*15/120 = $37,500
Year 2 19,333 $(300,000-39,990) * 13.33% = $34,659 $300,000*14/120 = $35,000
Year 3 19,333 $(300,000-39,990-34,659) *13.33% = $30,039 $300,000*13/120 = $32,500
Year 4 19,333 $(300,000-39,990-34,659-30,039)* 13.33% = $26,035 $300,000*12/120 = $30,000
Sum of year digit method: $300,000-10,000 * 12/120 $29,000

Doble declining balance method:

Depreciation rate = 100/15*2 = 13.33%

($300,000-13.33%) * 13.33% $34,659

Revised depreciation expense :

Depreciation expense for year1 and 2 = $300,000-10,000/15 = $19,333*2 = $38,666

Book value as on Year 3 = $300,000-38,666 = $261,334

$261,334-5,000/10 $25,633